Preparing a paper using L^{A}TEX 2ε for publication in Journal of Physics: Conference Series

Graham Douglas

Services Developer, Institute of Physics Publishing, Dirac House, Temple Back, Bristol BS1 6BE, UK

               E-mail: jpcs@iop.org

Abstract. This document describes the preparation of a conference paper to be published in Journal of Physics: Conference Series using L^{A}TEX 2ε and the jpconf.cls class file. This class file is designed to help authors produce their work in the best form for publication.

.

1. Introduction These guidelines show how to prepare articles for publication in Journal of Physics: Conference Series using L^{A}TEX so they can be published quickly and accurately. Articles will be refereed by the conference organizers but will not be edited or proof read after acceptence. It is, therefore, the author’s responsibility to ensure the accuracy of the content. This document has been prepared using jpconf.cls so serves as a sample document. The class file and accompanying documentation are available from http://jpconf.iop.org by downloading zipped files for PCs or tar compressed format files for Unix.

2. What you need to supply Authors must send their articles and files to the conference organizer, not direct to Institute of Physics Publishing (IOPP). Please pay particular attention to any specific instructions given to you by the conference organizer, especially with regard to the maximum length of the article. Appendix A contains a list of points to check before you supply anything to the conference organizer. In summary, you will need to provide:

• The L^{A}TEX source code, for the complete article with embedded figures, including any non-standard packages or macro files.

• Graphics files for all figures, as Encapsulated PostScript (EPS) files. For further guidance on preparing EPS files please refer to the graphics preparation guidelines which

    may be downloaded from http://jpconf.iop.org

• A PDF (or PostScript) file of the complete article.

• A printout of the article, to allow us to check what is required in the event of problems with the files.

• A signed Assignment of Copyright form. Available in PDF from http://jpconf.iop.org

Combine all your files into a single archive using a compression or archiving program prior to sending. Utilities like WinZIP or StuffIt will both archive and compress your files. Alternatively you could archive your files using a program like TAR and compress the resulting file using gzip.

3. Preparing your paper Authors may add their own macros at the start of an article provided they do not overwrite existing definitions and that they send copies of their new macros with their text file. jpconf requires L^{A}TEX 2ε and can be used with other package files such as those loading the AMS extension fonts msam and msbm (these fonts provide the blackboard bold alphabet and various extra maths symbols as well as symbols useful in figure captions); an extra style file iopams.sty is provided to load these packages and provide extra definitions for bold Greek letters.

3.1. Headers, footers and page numbers Authors should not add headers, footers or page numbers to the pages of their article—they will be added by Institute of Physics Publishing as part of the production process.

3.2. jpconf.cls package options The jpconf.cls class file has two options ‘a4paper’ and ‘letterpaper’:

\documentclass[a4paper]{jpconf}

or

\documentclass[letterpaper]{jpconf}

Table 1. jpconf.cls class file options.

Option  Description
a4paper  Set the paper size and margins for A4 paper.
letterpaper Set the paper size and margins for US letter paper.

The default paper size is A4 (i.e., the default option is a4paper) but this can be changed to

Letter by using \documentclass[letterpaper]{jpconf}.

4. The title, authors, addresses and abstract The code for setting the title page information is slightly different from the normal default in L^{A}TEX but please follow these instructions as carefully as possible so all articles within a conference have the same style to the title page. The title is set in bold unjustified type using the command \title{#1}, where #1 is the title of the article. The first letter of the title should be capitalized with the rest in lower case. The next information required is the list of all authors’ names followed by the affiliations. For the authors’ names type \author{#1}, where #1 is the list of all authors’ names. The style for the names is initials then surname, with a comma after all but the last two names, which are separated by ‘and’. Initials should not have full stops. First names may be used if desired. The command \maketitle is not required.

The addresses of the authors’ affiliations follow the list of authors. Each address should be set by using \address{#1} with the address as the single parameter in braces. If there is more than one address then a superscripted number, followed by a space, should come at the start of each address. In this case each author should also have a superscripted number or numbers following their name to indicate which address is the appropriate one for them.

Please also provide e-mail addresses for any or all of the authors using an \ead{#1} command after the last address. \ead{#1} provides the text Email: so #1 is just the e-mail address or a list of emails.

The abstract follows the addresses and should give readers concise information about the content of the article and should not normally exceed 200 words. To indicate the start of the abstract type \begin{abstract} followed by the text of the abstract. The abstract should normally be restricted to a single paragraph and is terminated by the command \end{abstract}

4.1. Sample coding for the start of an article The code for the start of a title page of a typical paper might read:

\title{The anomalous magnetic moment of the
neutrino and its relation to the solar neutrino problem}
\author{P J Smith$^1$, T M Collins$^2$,
R J Jones$^{3,}$\footnote[4]{Present address:
Department of Physics, University of Bristol, Tyndalls Park Road,
Bristol BS8 1TS, UK.} and Janet Williams$^3$}
\address{$^1$ Mathematics Faculty, Open University,
Milton Keynes MK7~6AA, UK}
\address{$^2$ Department of Mathematics,
Imperial College, Prince Consort Road, London SW7~2BZ, UK}
\address{$^3$ Department of Computer Science,
University College London, Gower Street, London WC1E~6BT, UK}
\ead{williams@ucl.ac.uk}
\begin{abstract}
The abstract appears here.
\end{abstract}

5. The text The text of the article should should be produced using standard L^{A}TEX formatting. Articles may be divided into sections and subsections, but the length limit provided by the conference organizer should be adhered to.

5.1. Acknowledgments Authors wishing to acknowledge assistance or encouragement from colleagues, special work by technical staff or financial support from organizations should do so in an unnumbered Acknowledgments section immediately following the last numbered section of the paper. The command \ack sets the acknowledgments heading as an unnumbered section.

5.2. Appendices Technical detail that it is necessary to include, but that interrupts the flow of the article, may be consigned to an appendix. Any appendices should be included at the end of the main text of the paper, after the acknowledgments section (if any) but before the reference list. If there are two or more appendices they will be called Appendix A, Appendix B, etc. Numbered equations will be in the form (A.1), (A.2), etc, figures will appear as figure A1, figure B1, etc and tables as table A1, table B1, etc.

The command \appendix is used to signify the start of the appendixes. Thereafter \section, \subsection, etc, will give headings appropriate for an appendix. To obtain a simple heading of ‘Appendix’ use the code \section*{Appendix}. If it contains numbered equations, figures or tables the command \appendix should precede it and \setcounter{section}{1} must follow it.

6. References In the online version of Journal of Physics: Conference Series references will be linked to their original source or to the article within a secondary service such as INSPEC or ChemPort wherever possible. To facilitate this linking extra care should be taken when preparing reference lists.

Two different styles of referencing are in common use: the Harvard alphabetical system and the Vancouver numerical system. For Journal of Physics: Conference Series, the Vancouver numerical system is preferred but authors should use the Harvard alphabetical system if they wish to do so. In the numerical system references are numbered sequentially throughout the text within square brackets, like this [2], and one number can be used to designate several references.

6.1. Reference lists A complete reference should provide the reader with enough information to locate the article concerned, whether published in print or electronic form, and should, depending on the type of reference, consist of:

• name(s) and initials;

• date published;

• title of journal, book or other publication;

• titles of journal articles may also be included (optional);

• volume number;

• editors, if any;

• town of publication and publisher in parentheses for books;

• the page numbers.

Up to ten authors may be given in a particular reference; where there are more than ten only the first should be given followed by ‘et al’. If an author is unsure of a particular journal’s abbreviated title it is best to leave the title in full. The terms loc. cit. and ibid. should not be used. Unpublished conferences and reports should generally not be included in the reference list and articles in the course of publication should be entered only if the journal of publication is known. A thesis submitted for a higher degree may be included in the reference list if it has not been superseded by a published paper and is available through a library; sufficient information should be given for it to be traced readily.

6.2. Formatting reference lists Numeric reference lists should contain the references within an unnumbered section (such as \section*{References}). The reference list itself is started by the code \begin{thebibliography}{<num>}, where <num> is the largest number in the reference list and is completed by \end{thebibliography}. Each reference starts with \bibitem{<label>}, where ‘label’ is the label used for cross-referencing. Each \bibitem should only contain a reference to a single article (or a single article and a preprint reference to the same article).

When one number actually covers a group of two or more references to different articles, \nonum should replace \bibitem{<label>} at the start of each reference in the group after the first.

  For an alphabetic reference list use \begin{thereferences} ... \end{thereferences}

instead of the ‘thebibliography’ environment and each reference can be start with just \item instead of \bibitem{label} as cross referencing is less useful for alphabetic references.

6.3. Using BibTEX For Journal of Physics: Conference Series, the standard unsrt.bst BibTEX style file is preferred. However, should you use any any other BibTEX .bst style file please be sure to include this, and your .bib file, with your submission to the conference organizer.

6.4. References to printed journal articles A normal reference to a journal article contains three changes of font (see table 2) and is constructed as follows:

• the authors should be in the form surname (with only the first letter capitalized) followed by the initials with no periods after the initials. Authors should be separated by a comma except for the last two which should be separated by ‘and’ with no comma preceding it;

• the article title (if given) should be in lower case letters, except for an initial capital, and should follow the date;

• the journal title is in italic and is abbreviated. If a journal has several parts denoted by different letters the part letter should be inserted after the journal in Roman type, e.g. Phys. Rev. A;

• the volume number should be in bold type;

• both the initial and final page numbers should be given where possible. The final page number should be in the shortest possible form and separated from the initial page number by an en rule ‘– ’, e.g. 1203–14, i.e. the numbers ‘12’ are not repeated.

A typical (numerical) reference list might begin

[1] Strite S and Morkoc H 1992 J. Vac. Sci. Technol. B 10 1237

[2] Jain S C, Willander M, Narayan J and van Overstraeten R 2000 J. Appl. Phys. 87 965

[3] Nakamura S, Senoh M, Nagahama S, Iwase N, Yamada T, Matsushita T, Kiyoku H and Sugimoto Y 1996 Japan. J. Appl. Phys. 35 L74

[4] Akasaki I, Sota S, Sakai H, Tanaka T, Koike M and Amano H 1996 Electron. Lett. 32 1105

[5] O’Leary S K, Foutz B E, Shur M S, Bhapkar U V and Eastman L F 1998 J. Appl. Phys. 83 826

[6] Jenkins D W and Dow J D 1989 Phys. Rev. B 39 3317

which would be obtained by typing

\begin{\thebibliography}{9}
\item Strite S and Morkoc H 1992 {\it J. Vac. Sci. Technol.} B {\bf 10} 1237
\item Jain S C, Willander M, Narayan J and van Overstraeten R 2000
{\it J. Appl. Phys}. {\bf 87} 965
\item Nakamura S, Senoh M, Nagahama S, Iwase N, Yamada T, Matsushita T, Kiyoku H
and  Sugimoto Y 1996 {\it Japan. J. Appl. Phys.} {\bf 35} L74
\item Akasaki I, Sota S, Sakai H, Tanaka T, Koike M and Amano H 1996
{\it Electron. Lett.} {\bf 32} 1105
\item O'Leary S K, Foutz B E, Shur M S, Bhapkar U V and Eastman L F 1998
{\it J. Appl. Phys.} {\bf 83} 826
\item Jenkins D W and Dow J D 1989 {\it Phys. Rev.} B {\bf 39} 3317
\end{\thebibliography}

Table 2. Font styles for a reference to a journal article.

Element  Style
Authors  Roman type
Date  Roman type
Article title (optional) Roman type
Journal title  Italic type
Volume number  Bold type
Page numbers  Roman type

Table 3. Font styles for references to books, conference proceedings and reports.

Element  Style
Authors  Roman type
Date  Roman type
Book title (optional)  Italic type
Editors  Roman type
Place (city, town etc) of publication Roman type
Publisher  Roman type
Volume  Roman type
Page numbers  Roman type

6.5. References to Journal of Physics: Conference Series articles Each conference proceeding published in Journal of Physics: Conference Series will be a separate volume; consequently references should follow the style for conventional printed journals. For example:

[1] Douglas G 2004 J. Phys.: Conf. Series 1 23–36

6.6. References to preprints For preprints there are two distinct cases:

(1) Where the article has been published in a journal and the preprint is supplementary reference information. In this case it should be presented as:

[1] Kunze K 2003 T-duality and Penrose limits of spatially homogeneous and inhomogeneous cosmologies

Phys. Rev. D 68 063517 (Preprint gr-qc/0303038)

(2) Where the only reference available is the preprint. In this case it should be presented as

[1] Milson R, Coley A, Pravda V and Pravdova A 2004 Alignment and algebraically special tensors Preprint gr-qc/0401010

6.7. References to books, conference proceedings and reports References to books, proceedings and reports are similar to journal references, but have only two changes of font (see table 3).

Points to note are:

• Book titles are in italic and should be spelt out in full with initial capital letters for all except minor words. Words such as Proceedings, Symposium, International, Conference, Second, etc should be abbreviated to Proc., Symp., Int., Conf., 2nd, respectively, but the rest of the title should be given in full, followed by the date of the conference and the town or city where the conference was held. For Laboratory Reports the Laboratory should be spelt out wherever possible, e.g. Argonne National Laboratory Report.

• The volume number, for example vol 2, should be followed by the editors, if any, in a form such as ‘ed A J Smith and P R Jones’. Use et al if there are more than two editors. Next comes the town of publication and publisher, within brackets and separated by a colon, and finally the page numbers preceded by p if only one number is given or pp if both the initial and final numbers are given.

Examples taken from published papers:

[1] Kurata M 1982 Numerical Analysis for Semiconductor Devices (Lexington, MA: Heath)

[2] Selberherr S 1984 Analysis and Simulation of Semiconductor Devices (Berlin: Springer)

[3] Sze S M 1969 Physics of Semiconductor Devices (New York: Wiley-Interscience)

[4] Dorman L I 1975 Variations of Galactic Cosmic Rays (Moscow: Moscow State University Press) p 103

[5] Caplar R and Kulisic P 1973 Proc. Int. Conf. on Nuclear Physics (Munich) vol 1 (Amsterdam: North-Holland/American Elsevier) p 517

[6] Cheng G X 2001 Raman and Brillouin Scattering-Principles and Applications (Beijing: Scientific)

[7] Szytula A and Leciejewicz J 1989 Handbook on the Physics and Chemistry of Rare Earths vol 12, ed K A Gschneidner Jr and L Erwin (Amsterdam: Elsevier) p 133

[8] Kuhn T 1998 Density matrix theory of coherent ultrafast dynamics Theory of Transport Properties of Semiconductor Nanostructures (Electronic Materials vol 4) ed E Schöll (London: Chapman and Hall) chapter 6 pp 173–214

7. Cross referencing L^{A}TEX provides excellent facilities for doing cross-referencing and these can be very useful in preparing articles. However, before you supply your files to the conference organizer make sure to check for any undefined references—which will generate a L^{A}TEX warning:

  LaTeX Warning:  There were undefined references.

Please check your printout and the L^{A}TEX log file for any undefined references before sending your files to the conference organizer.

8. Tables and table captions Tables should be numbered serially and referred to in the text by number (table 1, etc, rather than tab. 1). Each table should be a float and be positioned within the text at the most convenient place near to where it is first mentioned in the text. It should have an explanatory caption which should be as concise as possible.

8.1. The basic table format The standard form for a table is:

\begin{table}
\caption{\label{label}Table caption.}
\begin{center}
\begin{tabular}{llll}
\br
Head 1&Head 2&Head 3&Head 4\\
\mr
1.1&1.2&1.3&1.4\\
2.1&2.2&2.3&2.4\\
\br
\end{tabular}

\end{center}
\end{table}

The above code produces table 4.

Table 4. Table caption.

Head 1 Head 2 Head 3 Head 4
1.1  1.2  1.3  1.4
2.1  2.2  2.3  2.4

Points to note are:

(1) The caption comes before the table.

(2) The normal style is for tables to be centred in the same way as equations. This is

    accomplished by using \begin{center} . . . \end{center}.

(3) The default alignment of columns should be aligned left.

(4) Tables should have only horizontal rules and no vertical ones. The rules at the top and bottom are thicker than internal rules and are set with \br (bold rule). The rule separating the headings from the entries is set with \mr (medium rule). These commands do not need a following double backslash.

(5) Numbers in columns should be aligned as appropriate, usually on the decimal point; to help do this a control sequence \lineup has been defined which sets \0 equal to a space the size of a digit, \m to be a space the width of a minus sign, and \- to be a left overlapping minus sign. \- is for use in text mode while the other two commands may be used in maths or text. (\lineup should only be used within a table environment after the caption so that \- has its normal meaning elsewhere.) See table 5 for an example of a table where \lineup has been used.

Table 5. A simple example produced using the standard table commands and \lineup to assist in aligning columns on the decimal point. The width of the table and rules is set automatically by the preamble.

\begin{matrix}A & B & C & D & E & F & G \\ 23.5 & 60 & 0.53 & −20.2 & −0.22 & 1.7 & 14.5 \\ 39.7 & −60 & 0.74 & −51.9 & −0.208 & 47.2 & 146 \\ 123.7 & 0 & 0.75 & −57.2 & — & — & — \\ 3241.56 & 60 & 0.60 & −48.1 & −0.29 & 41 & 15 \end{matrix}

9. Figures and figure captions Figures must be included in the source code of an article at the appropriate place in the text not grouped together at the end. Figures should be included as encapsulated PostScript files using one of the standard L^{A}TEX graphics packages, (ideally graphicx or graphics) or using the L^{A}TEX picture environment.

Each figure should have a brief caption describing it and, if necessary, interpreting the various lines and symbols on the figure. As much lettering as possible should be removed from the figure itself and included in the caption. If a figure has parts, these should be labelled (a), (b), (c), etc. Table 6 gives the definitions for describing symbols and lines often used within figure captions (more symbols are available when using the optional packages loading the AMS extension fonts).

Table 6. Control sequences to describe lines and symbols in figure captions.

Control sequence Output  Control sequence  Output

        \dotted  · · · · · ·  \opencircle  ∘
             \dashed  - - - -  \opentriangle  △
             \broken  – – –  \opentriangledown ▽
             \longbroken  — — —  \fullsquare
             \chain  — · —  \opensquare  ⊓ ⊔
        \dashddot  — · · —  \fullcircle  •
             \full  ——  \opendiamond  ♢

Authors should try and use the space allocated to them as economically as possible. At times it may be convenient to put two figures side by side or the caption at the side of a figure. To put figures side by side, within a figure environment, put each figure and its caption into a minipage with an appropriate width (e.g. 3in or 18pc if the figures are of equal size) and then separate the figures slightly by adding some horizontal space between the two minipages (e.g. \hspace{.2in} or \hspace{1.5pc}. To get the caption at the side of the figure add the small horizontal space after the \includegraphics command and then put the \caption within a minipage of the appropriate width aligned bottom, i.e. \begin{minipage}[b]{3in} etc (see code in this file used to generate figures 1–3).

Note that it may be necessary to adjust the size of the figures (using optional arguments to \includegraphics, for instance [width=3in]) to get you article to fit within your page allowance or to obtain good page breaks.

Authors should ensure their EPS files meet the following criteria:

• The Bounding Box should indicate the area of the figure with a minimum of white space around it and not the dimensions of the page.

• Any fonts used should be from the standard PostScript set (Times, Helvetica, Courier and Symbol).

• Scanned images should be 600 dpi resolution for line art (black and white) and 300 dpi resolution for grayscale or colour.

• Captions and labels (e.g. Figure 1) should not be included in the EPS file although part letters (e.g. (a)) are acceptable provided they are placed close or within the boundary of the figure.

The precise coding required will depend on the graphics package being used and the printer driver. IOPP will use dvips but authors should avoid using special effects generated by including verbatim PostScript code within the L^{A}TEX file with specials other than the standard figure inclusion ones.

Figure 1. Figure caption for first of two sided figures.

Figure 2. Figure caption for second of two sided figures.

Figure 3. Figure caption for a narrow figure where the caption is put at the side of the figure.

Using the graphicx package figures can be included using code such as:

\begin{figure}
\begin{center}
\includegraphics{file.eps}
\end{center}
\caption{\label{label}Figure caption}
\end{figure}

Appendix A. Checklists Before you send any materials to the conference organizer, please read through the following checklist.

• Have you printed and signed an Institute of Physics Publishing Assignment of Copyright form?

• Have you prepared a printed copy of your paper and checked that it is free from printing errors—such as font problems, missing graphics etc?

• Does your manuscript have an abstract, which contains the essential information of the article? It should be complete by itself and suitable for direct inclusion in an abstracting service.

• Have you checked your references (including cross references):

– Have you performed two L^{A}TEX runs to check for unresolved references—i.e., any ‘??’ present in the printout?

– Does your manuscript have a complete list of references—preferably in the Vancouver (numerical) system?

– Have you done a literature search to check for relevant references you may have missed?

• Are all figures are numbered and referred to correctly in the text?

• Are all figures clear and readable?

• Are all acronyms and abbreviations are clearly explained when they first appear in the text, and the units used are consistent throughout the article?

• Do we have your current correspondence address as well as any address changes for the coming year?

Appendix A.1. Checking your files • If you have used any non-standard L^{A}TEX packages have you remembered to include them?

• If you have used BibTEX have you included the bibliographic database (.bib) file together with any (.bst) style files?

• Have you prepared/supplied all your graphics as EPS files?

• Have you included all the graphics in the L^{A}TEX 2ε source using the \includegraphics command?

• Have you prepared/supplied a PostScript or PDF file prepared from your paper?

• Have you checked the L^{A}TEX’s log file for any L^{A}TEX or package errors or warnings?